# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v3

@@ 1 Introduction
Subsampling algorithm. We model the problem as a Markov Decision Process (MDP) with a global agent and $n$ local agents. We propose SUBSAMPLE-MFQ to address the challenge of MARL with a large number of local agents. SUBSAMPLE-MFQ selects $k\leq n$ local agents to learn a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k}$ by applying mean-field value iteration on the $k$-local-agent subsystem to learn $\hat{Q}_{k}^{\mathrm{est}}$, which can be viewed as a smaller $Q$ function. SUBSAMPLE-MFQ then deploys a stochastic policy ${\pi}^{\mathrm{est}}_{k}$, where the global agent samples $k$ local agents uniformly at each step and uses $\hat{\pi}^{\mathrm{est}}_{k}$ to determine its action, while each local agent samples $k-1$ other local agents and uses $\hat{\pi}^{\mathrm{est}}_{k}$ to determine its action.
Sample complexity and theoretical guarantee. As the number of local agents increases, the size of $\hat{Q}_{k}$ scales polynomially with $k$, rather than polynomially with $n$ as in mean-field MARL. Analogously, when the size of the local agent’s state space grows, the size of $\hat{Q}_{k}$ scales exponentially with $k$, rather than exponentially with $n$, as in traditional $Q$-learning. The key analytic technique underlying our results is a novel MDP sampling result. Through it, we show that the performance gap between ${{\pi}_{k}^{\mathrm{est}}}$ and the optimal policy ${\pi^{*}}$ is at most $\tilde{O}(1/\sqrt{k})$, with high probability. The choice of $k$ reveals a fundamental trade-off between the size of the $Q$-table and the optimality of ${\pi}_{k}^{\mathrm{est}}$. For example, if $k$ is set to $O(\log n)$, SUBSAMPLE-MFQ is the first centralized MARL algorithm to achieve a polylogarithmic run-time in $n$, representing an exponential speedup over the previously best-known polytime mean-field MARL methods, while maintaining a decaying optimality gap as $n$ gets large,
While our results are theoretical in nature, we hope SUBSAMPLE-MFQ will further exploration of sampling in Markov games, and inspire new practical multi-agent algorithms.
@@
Other related works. Our work is related to factored MDPs, where there is a global action affecting every agent; however, in our case, each agent has its own action (Min et al., 2023; Lauer and Riedmiller, 2000). Jin et al. (2020) reduces the dependence of the product action space to an additive dependence with V-learning. Our work further reduces the complexity of the joint state space, which has not been previously accomplished. We add to the growing literature on the Centralized Training with Decentralized Execution regime (Zhou et al., 2023), as our algorithm learns a provably approximately optimal policy using centralized information, but makes decisions using only local information during execution. Finally, one can efficiently approximate the $Q$-table through function approximation (Jin et al., 2021). However, achieving theoretical bounds on the performance loss due to function approximation is intractable without strong assumptions such as linear Bellman completeness or low Bellman-Eluder dimension (Golowich and Moitra, 2024). While our work primarily studies the finite tabular setting, we extend it to non-tabular linear MDPs in Appendix J.
@@ 2 Preliminaries
For $k,n\in\mathbb{N}$ where $k\leq n$, let $\binom{[n]}{k}$ denote the set of $k$-sized subsets of $[n]=\{1,\dots,n\}$. For any vector $z\in\mathbb{R}^{d}$, let $\|z\|_{1}$ and $\|z\|_{\infty}$ denote the standard $\ell_{1}$ and $\ell_{\infty}$ norms of $z$ respectively. Let $\|\mathbf{A}\|_{1}$ denote the matrix $\ell_{1}$-norm of $\mathbf{A}\in\mathbb{R}^{n\times m}$. Given variables $s_{1},\dots,s_{n}$, $s_{\Delta}\coloneqq\{s_{i}:i\in\Delta\}$ for $\Delta\subseteq[n]$. We use $\tilde{O}(\cdot)$ to suppress polylogarithmic factors in all problem parameters except $n$. For a discrete measurable space $(\mathcal{X},\mathcal{F})$, the total variation distance between probability measures $\rho_{1},\rho_{2}$ is given by $\mathrm{TV}(\rho_{1},\rho_{2})=\frac{1}{2}\sum_{x\in\mathcal{X}}|\rho_{1}(x)-\rho_{2}(x)|$. Next, $x\sim\mathcal{D}(\cdot)$ denotes that $x$ is a random element sampled from a distribution $\mathcal{D}$, and we denote that $x$ is a random sample from the uniform distribution over a finite set $\Omega$ by $x\sim\mathcal{U}(\Omega)$. We include a detailed notation table in Table 1.
@@ 2.1 Problem formulation
The system then collects a structured stage reward $r(s(t),a(t))$ where the reward $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ depends on $s(t)$ and $a(t)$ through Equation 2, and where $r_{g}$ and $r_{l}$ is typically application specific.
@@
We assume that the state and action spaces of all the agents in the MARL game are finite: $|\mathcal{S}_{l}|,|\mathcal{S}_{g}|,|\mathcal{A}_{g}|,|\mathcal{A}_{l}|<\infty$. Appendix J of the supplementary material relaxes this assumption to the non-tabular setting with infinite continuous sets.
@@
Motivating examples. Below we give examples of two cooperative MARL settings which are naturally modeled by our setting. Our experiments in Appendix B reveal a monotonic improvement in the learned policies as $k\to n$, while providing a substantial speedup over mean-field $Q$-learning.
@@
  • Constrained exploration: Consider an $M\times M$ grid. Each agent’s state is a coordinate in $[M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent constrains the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns. Through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while minimizing risk.
@@
Following Mondal et al. (2022); Xu and Klabjan (2023), our model can capture heterogeneity in the local agents by modeling agent types as part of the state: assign a type $\varepsilon_{i}\in\mathcal{E}$ to each local agent by letting $\mathcal{S}_{l}=\mathcal{E}\times\mathcal{S}^{\prime}_{l}$, where $\mathcal{E}$ is a set of possible types that are treated as a fixed part of the agent’s state. The transition and reward functions can vary depending on the agent’s type. The global agent can provide unique signals to local agents of each type by letting $s_{g}\in\mathcal{S}_{g}$ and $a_{g}\in\mathcal{A}_{g}$ denote a state/action vector where each element corresponds to a type.
@@ 3 Algorithmic Approach: Subsampled Value Iteration
Offline Planning: Algorithm 1. First, the global agent randomly samples a subset of local agents $\Delta\subseteq[n]$ such that $|\Delta|=k$, for $k\leq n$. It then ignores all other local agents $[n]\setminus\Delta$, and performs value iteration (using $m$ samples to update the $Q$-function in each iterate) to approximately learn the $Q$-function $\hat{Q}_{k,m}^{\mathrm{est}}$ and policy $\hat{\pi}_{k,m}^{\mathrm{est}}$ for this surrogate subsystem of $k$ local agents. We denote the surrogate reward gained by this subsystem at each time step by $r_{\Delta}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, where
@@
In Theorem E.3, we show that $\|\hat{Q}_{k,m}^{\mathrm{est}}-Q^{*}\|_{\infty}$ is Lipschitz continuous with respect to the TV-distance between the state/action pairs of the subsampled agents and the full set of $n$ agents. Equipped with this approximation guarantee, we show how to construct an approximately optimal policy on the full system on $n$ agents. In general, converting this policy on $k$ local agents to a policy on the full $n$-agent system without sacrificing error guarantees can be intractable, and there is a line on works on centralized-training decentralized-execution (CTDE) (Xu and Klabjan, 2023; Zhou et al., 2023) which shows that designing performant decentralized policies can be highly non-trivial.
Online Execution: Algorithm 2. In order to circumvent this obstacle and convert the optimality of each agent’s action in the $k$ local-agent subsystem to an approximate optimality guarantee on the full $n$-agent system, we propose a randomized policy ${\pi}^{\mathrm{est}}_{k,m}$, where the global agent samples $\smash{\Delta\in{[n]\choose k}}$ at each time-step to derive an action $\smash{a_{g}\leftarrow\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})}$, and where each local $i$ agent samples $\smash{k-1}$ other local agents $\Delta_{i}$ to derive an action $\smash{a_{i}\leftarrow\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta_{i}})}$. Finally, Theorem 4.4 shows that the policy ${\pi}_{k,m}^{\mathrm{est}}$ converges to the optimal policy $\pi^{*}$ as $\smash{k\to n}$ with rate $\tilde{O}(\frac{1}{\sqrt{k}})$.
@@
We now formally describe the algorithms. Before describing Algorithm 1 (SUBSAMPLE-MFQ: Learning) and Algorithm 2 (SUBSAMPLE-MFQ: Execution) in detail, it will be helpful to first define the empirical distribution function:
@@
Algorithm 1 (Offline learning). Let $m\in\mathbb{N}$ denote the sample size for the learning algorithm with sampling parameter $k\leq n$. As in Remark 3.1, when $\smash{|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}}$, the algorithm uses traditional value-iteration, and when $\smash{|\mathcal{Z}_{l}|^{k}>|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}}$, it uses mean-field value iteration. We formally describe the procedure for each regime below:
Regime with Large State/Action Space: When $|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$, we iteratively learn the optimal $Q$-function for a subsystem with $k$-local agents denoted by $\smash{\hat{Q}_{k,m}^{t}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}\to\mathbb{R}}$, which is initialized to 0. At time $t$, we update $\hat{Q}_{k,m}^{t+1}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})={\tilde{\mathcal{T}}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, where ${\tilde{\mathcal{T}}}_{k,m}$ is the empirically adapted Bellman operator in Equation 7. Since the system is unknown, $\tilde{\mathcal{T}}_{k,m}$ cannot directly perform the Bellman update, so it instead uses $m$ random samples $s_{g}^{j}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\sim P_{l}(\cdot|s_{i},s_{g},a_{i})$ for each $j\in[m]$, $i\in\Delta$ to approximate the system:
${\tilde{\mathcal{T}}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k,m}^{t}(s_{g}^{j},{s_{\Delta}^{j}},a_{g}^{\prime},a_{\Delta}^{\prime}).$ (7)
Since $\tilde{\mathcal{T}}_{k,m}$ is $\gamma$-contractive, Algorithm 1 applies value iteration with $\tilde{\mathcal{T}}$ until $\hat{Q}_{k,m}$ converges to a fixed point satisfying $\smash{\tilde{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{\mathrm{est}}=\hat{Q}_{k,m}^{\mathrm{est}}}$, yielding a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ where
$\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ $\!=\!\!\mathop{\arg\max}_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}},a_{g},a_{\Delta}).$
@@
0:  A multi-agent system as in Section 2, number of iterations $T$, sampling parameters $k\in[n]$, $m\in\mathbb{N}$, and discount factor $\gamma\in(0,1)$.
@@
Regime with Large Number of Agents: When $|\mathcal{Z}_{l}|^{k}>|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$, we learn the optimal mean-field $Q$-function for a $k$ local agent system, denoted by $\hat{Q}_{k,m}^{t}:\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}\to\mathbb{R}$, which is initialized to 0. At time $t$, we update $\hat{Q}_{k,m}^{t+1}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})=\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})$, where $\hat{\mathcal{T}}_{k,m}$ is the empirically adapted mean-field Bellman operator in Equation 8. Similarly, as the system is unknown, $\hat{\mathcal{T}}_{k,m}$ cannot directly perform the Bellman update and instead uses $m$ random samples $s_{g}^{j}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall j\in[m]$, $i\in\Delta$ to approximate the system:
@@
Since $\hat{Q}_{k,m}^{t}$ depends on $s_{\Delta}$ and $a_{\Delta}$ through $F_{z_{\Delta}}$, $\hat{\mathcal{T}}_{k,m}$ is also $\gamma$-contractive. So, Algorithm 1 applies value iteration with $\hat{\mathcal{T}}$ until $\hat{Q}_{k,m}$ converges to a fixed point satisfying $\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{\mathrm{est}}=\hat{Q}_{k,m}^{\mathrm{est}}$, yielding a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\tilde{\Delta}}})$:
$\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\tilde{\Delta}}})$ $\!=\!\!\mathop{\arg\max}_{\begin{subarray}{c}a_{g}\in\mathcal{A}_{g},a_{1}\in\mathcal{A}_{l},F_{a_{\tilde{\Delta}}}\in\mu_{k-1}(\mathcal{A}_{l})\end{subarray}}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})$
For $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})\!=\!a_{g}^{*},a_{i}^{*},a_{\Delta\setminus i}^{*}$, let $[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}\!=\!a_{g}^{*}$ and $[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}\!=\!a_{i}^{*}$.
@@
Since mean-field $Q$-learning maintains the same updates as standard $Q$-learning (from Lemma C.19 which follows by noting that the $Q$-function is permutation-invariant in the homogeneous local agents), the deterministic policies $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})$ and $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\bar{\Delta}}})$ are equivalent.
Algorithm 2 (Online implementation). In Algorithm 2, (SUBSAMPLE-MFQ: Execution) the global agent samples local agents $\Delta(t)\sim\mathcal{U}\binom{[n]}{k}$ at each step to derive action $a_{g}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta}(t))]_{g}$, and each local agent $i$ samples other local agents $\Delta_{i}(t)\sim\mathcal{U}{[n]\setminus i\choose k-1}$ to derive action $a_{i}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta}(t))]_{l}$. The system then incurs a reward $r(s,a)$. This procedure of first sampling agents and then applying $\hat{\pi}^{\mathrm{est}}_{k,m}$ is denoted by a stochastic policy ${\pi}_{k,m}^{\mathrm{est}}(a|s)$, where ${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)$ is the global agent’s action distribution and ${\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)$ is the local agent’s action distribution:
@@
1:  Learn $\hat{\pi}_{k,m}^{\mathrm{est}}$ from Algorithm 1.
@@
6:     Choose $\Delta$ uniformly at random from (${[n]\choose k}$ and let $a_{g}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta}(t))]_{g}$.

### v4

@@ 1 Introduction
Subsampling algorithm. We model the problem as a Markov Decision Process (MDP) with a global agent and $n$ local agents. We propose SUBSAMPLE-MFQ to address the challenge of MARL with a large number of local agents. SUBSAMPLE-MFQ selects $k\leq n$ local agents to learn a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k}$ by applying mean-field value iteration on the $k$-local-agent subsystem to learn $\hat{Q}_{k}^{\mathrm{est}}$, which can be viewed as a smaller $Q$ function. SUBSAMPLE-MFQ then deploys a stochastic policy ${\pi}^{\mathrm{est}}_{k}$, where the global agent samples $k$ local agents uniformly at each step and uses $\hat{\pi}^{\mathrm{est}}_{k}$ to determine its action, while each local agent samples k1k-1 other local agents and uses $\hat{\pi}^{\mathrm{est}}_{k}$ to determine its action.
Sample complexity and theoretical guarantee. As the number of local agents increases, the size of $\hat{Q}_{k}$ scales polynomially with $k$, rather than polynomially with $n$ as in mean-field MARL. Analogously, when the size of the local agent’s state space grows, the size of $\hat{Q}_{k}$ scales exponentially with $k$, rather than exponentially with $n$, as in traditional $Q$-learning. The key analytic technique underlying our results is a novel MDP sampling result. This result shows that the performance gap between ${\pi}_{k}^{\mathrm{est}}$} and the optimal policy $\pi^{*}$} is at most $\tilde{O}(1/\sqrt{k})$, with high probability. The choice of $k$ reveals a fundamental trade-off between the size of the $Q$-table and the optimality of ${\pi}_{k}^{\mathrm{est}}$. For example, if $k$ is set to $O(\log n)$, SUBSAMPLE-MFQ is the first centralized MARL algorithm to achieve a polylogarithmic run-time in $n$, representing an exponential speedup over the previously best-known polytime mean-field MARL methods, while maintaining a decaying optimality gap as $n$ gets large,
While our results are theoretical in nature, we hope SUBSAMPLE-MFQ will further exploration of sampling in Markov games, and potentially inspire new practical multi-agent algorithms.
@@
Other related works. Our work is related to factored MDPs, where there is a global action affecting every agent; however, in our case, each agent has its own action (Min et al., 2023; Lauer and Riedmiller, 2000). Jin et al. (2020) reduces the dependence of the product action space to an additive dependence with V-learning. Our work further reduces the complexity of the joint state space, which has not been previously accomplished. We add to the growing literature on the Centralized Training with Decentralized Execution regime (Zhou et al., 2023), as our algorithm learns a provably approximately optimal policy using centralized information, but makes decisions using only local information during execution. Finally, one can efficiently approximate the $Q$-table through function approximation (Jin et al., 2021). However, achieving theoretical bounds on the performance loss due to function approximation is intractable without strong assumptions such as linear Bellman completeness or low Bellman-Eluder dimension (Golowich and Moitra, 2024). While our work primarily studies the finite tabular setting, we extend it to non-tabular linear MDPs in Appendix˜J.
@@ 2 Preliminaries
[…]
@@ 2.1 Problem formulation
The system then collects a structured stage reward $r(s(t),a(t))$ where the reward $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ depends on $s(t)$ and $a(t)$ through Equation˜2, and where $r_{g}$ and $r_{l}$ is typically application specific.
@@
We assume that the state and action spaces of all the agents in the MARL game are finite: $|\mathcal{S}_{l}|,|\mathcal{S}_{g}|,|\mathcal{A}_{g}|,|\mathcal{A}_{l}|<\infty$. Appendix˜J of the supplementary material relaxes this assumption to the non-tabular setting with infinite continuous sets.
@@
Motivating examples. Below we give examples of two cooperative MARL settings which are naturally modeled by our setting. Our experiments in Appendix˜B reveal a monotonic improvement in the learned policies as $k\to n$, while providing a substantial speedup over mean-field $Q$-learning [^1] We provide supporting code for the algorithm and experiments in https://github.com/emiletimothy/Mean-Field-Subsample-Q-Learning.
@@
  • Constrained exploration: Consider an $M\times M$ grid. Each agent’s state is a coordinate in $M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent constrains the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns, causing exploration in these regions to be disincentivized from the local agents’ perspectives. Through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while minimizing risk.
@@
Following Mondal et al. (2022); Xu and Klabjan (2023), our model can capture heterogeneity in the local agents by modeling agent types as part of the state: to do this, we assign a type $\varepsilon_{i}\in\mathcal{E}$ to each local agent by letting $\mathcal{S}_{l}=\mathcal{E}\times\mathcal{S}^{\prime}_{l}$, where $\mathcal{E}$ is a set that enumerates possible types that are treated as a fixed part of the agent’s state, and $\mathcal{S}_{l}^{\prime}$ is the latent state space of any local agent. The transition and reward functions can vary depending on the agent’s type. The global agent can provide unique signals to local agents of each type by letting $s_{g}\in\mathcal{S}_{g}$ and $a_{g}\in\mathcal{A}_{g}$ denote a state/action vector where each element corresponds to a type.
@@ 3 Algorithmic Approach: Subsampled Value Iteration
Offline Planning: Algorithm˜1. First, the global agent randomly samples a subset of local agents $\Delta\subseteq[n]$ such that $|\Delta|=k$, for $k\leq n$. It then ignores all other local agents $n]\setminus\Delta$, and performs value iteration (using $m$ samples to update the $Q$-function in each iterate) to approximately learn the $Q$-function $\hat{Q}_{k,m}^{\mathrm{est}}$ and policy $\hat{\pi}_{k,m}^{\mathrm{est}}$ for this surrogate subsystem of $k$ local agents. We denote the surrogate reward gained by this subsystem at each time step by $r_{\Delta}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, where
@@
In Theorem˜E.3, we show that $\|\hat{Q}_{k,m}^{\mathrm{est}}-Q^{*}\|_{\infty}$ is Lipschitz continuous with respect to the TV-distance between the state/action pairs of the subsampled agents and the full set of $n$ agents. Equipped with this approximation guarantee, we show how to construct an approximately optimal policy on the full system on $n$ agents. In general, converting this policy on $k$ local agents to a policy on the full $n$-agent system without sacrificing error guarantees can be intractable, and there is a line on works on centralized-training decentralized-execution (CTDE) (Xu and Klabjan, 2023; Zhou et al., 2023) which shows that designing performant decentralized policies can be highly non-trivial.
Online Execution: Algorithm˜2. In order to circumvent this obstacle and convert the optimality of each agent’s action in the $k$ local-agent subsystem to an approximate optimality guarantee on the full $n$-agent system, we propose a randomized policy ${\pi}^{\mathrm{est}}_{k,m}$, where the global agent samples $\smash{\Delta\in{[n]\choose k}}$ at each time-step to derive an action $\smash{a_{g}\leftarrow\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})}$, and where each local $i$ agent samples $\smash{k-1}$ other local agents $\Delta_{i}$ to derive an action $\smash{a_{i}\leftarrow\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta_{i}})}$. Finally, Theorem˜4.4 shows that the policy ${\pi}_{k,m}^{\mathrm{est}}$ converges to the optimal policy $\pi^{*}$ as $\smash{k\to n}$ with rate $\tilde{O}(\frac{1}{\sqrt{k}})$.
@@
We now formally describe the algorithms. Before describing Algorithm˜1 (SUBSAMPLE-MFQ: Learning) and Algorithm˜2 (SUBSAMPLE-MFQ: Execution) in detail, it will be helpful to first define the empirical distribution function:
@@
Algorithm˜1 (Offline learning). Let $m\in\mathbb{N}$ denote the sample size for the learning algorithm with sampling parameter $k\leq n$. As in Remark˜3.1, when $\smash{|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}}$, the algorithm uses traditional value-iteration, and when $\smash{|\mathcal{Z}_{l}|^{k}>|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}}$, it uses mean-field value iteration. We formally describe the procedure for each regime below:
Regime with Large State/Action Space: When $|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$, we iteratively learn the optimal $Q$-function for a subsystem with $k$-local agents denoted by $\smash{\hat{Q}_{k,m}^{t}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}\to\mathbb{R}}$, which is initialized to 0. At time $t$, we update
$\hat{Q}_{k,m}^{t+1}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})={\tilde{\mathcal{T}}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{\Delta},a_{g},a_{\Delta}),$ (7)
where ${\tilde{\mathcal{T}}}_{k,m}$ is the empirically adapted Bellman operator in Section˜3.2. Since the system is unknown, $\tilde{\mathcal{T}}_{k,m}$ cannot directly perform the Bellman update, so it instead uses $m$ random samples $s_{g}^{j}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\sim P_{l}(\cdot|s_{i},s_{g},a_{i})$ for each $j\in[m]$, $i\in\Delta$ to approximate the system:
${\tilde{\mathcal{T}}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k,m}^{t}(s_{g}^{j},{s_{\Delta}^{j}},a_{g}^{\prime},a_{\Delta}^{\prime}).$
Since $\tilde{\mathcal{T}}_{k,m}$ is $\gamma$-contractive, Algorithm˜1 applies value iteration with $\tilde{\mathcal{T}}$ until $\hat{Q}_{k,m}$ converges to a fixed point satisfying $\smash{\tilde{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{\mathrm{est}}=\hat{Q}_{k,m}^{\mathrm{est}}}$, yielding a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ where
$\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ $\!=\!\!\mathop{\arg\max}_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}},a_{g},a_{\Delta}).$ (8)
@@
0: A multi-agent system as in Section˜2, number of iterations $T$, sampling parameters $k\in[n]$, $m\in\mathbb{N}$, and discount factor $\gamma\in(0,1)$.
@@
Regime with Large Number of Agents: When $|\mathcal{Z}_{l}|^{k}>|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$, we learn the optimal mean-field $Q$-function for a $k$ local agent system, denoted by $\hat{Q}_{k,m}^{t}:\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}\to\mathbb{R}$, which is initialized to 0. At time $t$, we update
$\hat{Q}_{k,m}^{t+1}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})=\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g}),$ (9)
where $\hat{\mathcal{T}}_{k,m}$ is the empirically adapted mean-field Bellman operator in Equation˜10. Similarly, as the system is unknown, $\hat{\mathcal{T}}_{k,m}$ cannot directly perform the Bellman update and instead uses $m$ random samples $s_{g}^{j}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall j\in[m]$, $i\in\Delta$ to approximate the system:
@@
[…]
For $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})\!=\!a_{g}^{*},a_{i}^{*},a_{\Delta\setminus i}^{*}$, let $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}\!=\!a_{g}^{*}$ and $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}\!=\!a_{i}^{*}$.
@@
Since mean-field $Q$-learning maintains the same updates as standard $Q$-learning (from Lemma˜C.19 which follows by noting that the $Q$-function is permutation-invariant in the homogeneous local agents), the deterministic policies $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})$ and $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\bar{\Delta}}})$ are equivalent.
Algorithm˜2 (Online implementation). In Algorithm˜2, (SUBSAMPLE-MFQ: Execution) the global agent samples local agents $\Delta(t)\sim\mathcal{U}\binom{[n]}{k}$ at each step to derive action $a_{g}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta}(t))]_{g}$, and each local agent $i$ samples other local agents $\Delta_{i}(t)\sim\mathcal{U}{[n]\setminus i\choose k-1}$ to derive action $a_{i}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta}(t))]_{l}$. The system then incurs a reward $r(s,a)$. This procedure of first sampling agents and then applying $\hat{\pi}^{\mathrm{est}}_{k,m}$ is denoted by a stochastic policy ${\pi}_{k,m}^{\mathrm{est}}(a|s)$, where ${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)$ is the global agent’s action distribution and ${\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)$ is the local agent’s action distribution:
@@
1: Learn $\hat{\pi}_{k,m}^{\mathrm{est}}$ from Algorithm˜1.
@@
6:  Choose $\Delta$ uniformly at random from $n]\choose k$} and let $a_{g}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta}(t))]_{g}$.